Given the following normal distribution what is the probability X>35. That is, P(X>35). Assume standard deviation σ=6.
1
Expert's answer
2020-05-06T18:25:04-0400

If "X\sim N(\mu, \sigma^2)," then


"Z={X-\mu\over \sigma}\sim N(0, 1)"

Given "\mu=37, \sigma=6."


"P(X>35)=1-P(X\leq35)="

"=1-P(Z\leq{35-37\over 6})\approx1-P(Z\leq-0.333333)\approx"


"\approx1-0.369441=0.630559"